OpenStudy (anonymous):

Find f'(a) f(t) = 2t^3 + t

OpenStudy (anonymous):

Use \[ f(t) = a t^n + b t\\ f'(t) = n a t^{n-1} +b \]

OpenStudy (anonymous):

In your case, \[ f'(t) = 6 t^2 +1 \]

OpenStudy (anonymous):

hmm, we were taught to use f(a-h) - f(a)/h

OpenStudy (waleed_imtiaz):

So u are asking to derivate the given function.... So just use the power formula..... it will be 6t^2+1

OpenStudy (zzr0ck3r):

you sure it was not (f(a+h)-f(a))/h?

OpenStudy (waleed_imtiaz):

and if u want to use f(a-h) - f(a)/h or (f(a+h)-f(a))/h.... just put t=a+h or t=a-h...... and then t=a... and put in the formula.........

OpenStudy (zzr0ck3r):

huge difference on f(a-h) and f(a+h)

OpenStudy (zzr0ck3r):

(2(a+h)^3+(a+h)-(2a^3-h))/h = (2(a+h)(a^2+2ah+h^2)-2a^3+h))/h = 2(a^3+2a^2h+ah^2+ha^2+2ah^2+h^3)-2a^3+h)/h = (2a^3+4a^2h+2ah^2+2ha^2+4ah^2+2h^3-2a^3+h)/h = (6a^2h+6ah^2+2h^3+h)/h = 6a^2 + 6ah + 2h^2 + 1 run the limit can you do the next step?

OpenStudy (zzr0ck3r):

@dfresenius ?

OpenStudy (anonymous):

oh sorry was away,

OpenStudy (anonymous):

well I ended up with 6a^2+1 after running the limit.

OpenStudy (anonymous):

i was away working out all of those steps hadn't realized you wrote them

OpenStudy (zzr0ck3r):

np m8, you are right

OpenStudy (zzr0ck3r):

make sure to replace your a with a t so we have 6t^2+1

OpenStudy (anonymous):

You should have phrased your question by saying find the derivative using the definition as the limit of \[ \lim_{h->0} \frac {f(x+h)-f(x)}{h} \]